# Deep Legendre Transform

### v1

@@ 5.1 Numerical approximations with MLPs, ResNets, ICNNs
Our experiments with several convex functions across different dimensions [^8] Additional multidimensional experiments demonstrating that DLT is statistically identical to direct learning of the known conjugate in terms of time and errors are in the supplementary material. demonstrate that the performance of DLT is identical to direct learning in terms of both residual errors and computation time, as shown in Table 2. This equivalence is maintained across all network architectures tested, confirming that our method’s performance is only constrained by the neural networks’ inherent ability to approximate multidimensional functions, not by the implicit formulation itself. ResNet consistently provides the best approximation results despite not guaranteeing convexity, while MLP_ICNN shows limitations due to the lack of residual connections, resulting in significantly higher error rates across all test functions.
Table 2: Benchmark comparison of DLT vs. direct learning (Dir.) for $d=10$ and $50$. Errors are computed via Monte-Carlo over $4096$ i.i.d. test samples from $\text{Unif}\circ(\nabla f)^{-1}$ distribution.
@@
$_{\text{DLT}} / RMSE$$_{\text{Dir}}$ tsolve (s) $_{\text{DLT}} / RMSE$$_{\text{Dir}}$ tsolve (s)
@@
Functions without closed form conjugate.
Most functions have no closed-form expression for the convex conjugate. However, DLT gives the same posterior error estimates as if $f^{*}$ is known in a closed form and learned directly on $D$. For illustration, we consider the quadratic over linear function
$f(x)=\frac{x^{2}_{1}+\dots+x^{2}_{d}+1}{x_{1}+\dots+x_{d}+1}.$ (5.1)
It is not defined everywhere on $\mathbb{R}^{d}$. But it can be shown that it is convex on the convex set $C=\{x\in\mathbb{R}^{d}:x_{1}+\dots+x_{d}>0\}$; see e.g. boyd2004convex. The $i$-th component of the gradient $\nabla f(x)$ is
@@
But the convex conjugate $f^{*}$ is not known in closed form, and the image $D=\nabla f(C)$ is a non-trivial subset of $\mathbb{R}^{d}$. Even though we do not know $f^{*}$ explicitly, our method yields accurate ICNN approximations of $f^{*}$ on $D$. Here we used an ICNN with 3 hidden layers and ELU activations, we sampled random points from a standard half-normal distribution on $C$. Figure 1 depicts these random points and their images under the gradient map $\nabla f$ given in (5.2) for $d=2$. The table on the right shows estimates of the $L^{2}$-approximation error $\left\|g_{\hat{\theta}}-f^{*}\right\|_{L^{2}(\mathbb{R}^{d},\nu)}$ together with the times in minutes and seconds it took to train the approximations  $g_{\hat{\theta}}$ for different $d$. Note that the true $f^{*}$ is unknown here, but due to the results of Section 3, the DLT error is identical to the explicitly evaluated $L^{2}$-approximation error.
In Appendix B, we present additional non-separable examples: (i) quadratic $f(x)=\frac{1}{2}x^{\top}Qx$; (ii) exponential-minus-linear $f(x)=e^{\langle a,x\rangle}-\langle b,x\rangle$; and (iii) ICNN-represented functions.
$d$ $\|g_{\theta}-f^{*}\|_{L^{2}}$ Time
@@ 5.2 Comparison with classical grid methods
Here, we compare DLT against Lucet’s LLT algorithm (Lucet1997), a state-of-the-art grid-based method for computing convex conjugates. Rather than directly evaluating $\max_{x\in\mathcal{X}}\{\langle x,y\rangle-f(x)\}$ over a full $d$-dimensional grid $\mathcal{X}$ of size $N^{d}$ (which requires $O(N^{2d})$ operations), Lucet reformulates the $d$-dimensional maximization as a sequence of nested 1D problems:
$f^{*}(y_{1},\ldots,y_{d})=\max_{x_{1}}\left\{y_{1}x_{1}+\max_{x_{2}}\left\{y_{2}x_{2}+\cdots+\max_{x_{d}}\left\{y_{d}x_{d}-f(x)\right\}\cdots\right\}\right\}.$
This decomposition reduces computational complexity to $O(dN^{d+1})$ while still requiring $O(N^{d})$ memory. Table 3 compares the methods at a fixed discretization of $N=10$ points per dimension for the test setup from Table 1. All errors are evaluated on the same random sample $\mathcal{Y}_{\rm test}\subset D=\nabla f(C)$ against the true $f^{*}$. For low dimensions ($d\in\{2,6\}$), Lucet runs quickly with modest memory, making it feasible to use much finer grids within a reasonable time budget. However, as $d$ grows beyond 7788, even modest grids become memory or time infeasible due to $O(N^{d})$ storage and $O(d\cdot N^{d+1})$ computational complexity. In contrast, DLT maintains a constant memory footprint and scales smoothly to higher dimensions ($d\geq 9$).
Table 3: Comparison of Lucet with DLT across dimensions $d\in\{2,6,8,10\}$. Lucet uses $N=10$ points per dim; DLT uses a ResNet(128,128)(128,128). Reported: solution time $t_{\mathrm{solve}}$, evaluation time $t_{\mathrm{eval}}$, active memory (MB), and RMSE.
@@
When the gradient mapping $\nabla f$ is significantly distorting, the image $\nabla f(\mathcal{X})$ of a uniformly sampled training set $\mathcal{X}\subset C$ will be highly non-uniformly distributed on $D$. This skewed distribution can adversely affect learning dynamics, concentrating approximation accuracy in certain regions while neglecting others. The approximate inverse sampling method introduced in Section 4 addresses this issue by allowing us to sample from any desired distribution (e.g., uniform) directly on $D$.
@@
Table 4: Comparison of direct and inverse sampling for the negative logarithm. Inverse sampling adds an inverse map $h_{\hat{v}}$ with quality measured by $2/(s\sqrt{d})\|\nabla f\circ h_{\hat{v}}-\text{id}\|_{L^{1}}$, $s=990$ is the side-length of $D$, and total training time split between $h_{\hat{v}}$ and $g_{\hat{\theta}}$.
@@ 5.4 Applications to Hamilton–Jacobi equations
The Hamilton–Jacobi (HJ) equation is a cornerstone of classical mechanics, optimal control theory, and mathematical physics. It provides a natural application domain for our DLT methodology.
@@
where $H:\mathbb{R}^{d}\to\mathbb{R}$ is convex, and $g:\mathbb{R}^{d}\to\mathbb{R}$ is strictly convex and coercive.
The Hopf formula, hopf1965generalized, provides an analytical expression for the solution:
@@
which expresses the solution as a double convex conjugate. While this formula is explicit, evaluating it in high dimensions remains computationally demanding.
To address this, we propose a modified approach: the Time-Parameterized Deep Legendre Transform (Time-DLT): method approximates the function $(g^{*}+tH)^{*}$ directly using a neural network $g_{\theta}(y,t)$ that satisfies:
$g_{\theta}(\nabla u(x,t),t)\approx\langle x,\nabla u(x,t)\rangle-u(x,t),$ (5.3)
with $t$ appearing as a parameter. In the experimental setup the primal domain is $C=(-a,a)^{d}\times[0,T)$.
@@
We examine the case of a quadratic initial condition with a quadratic Hamiltonian  [^9] Here the inverse gradient mapping $h(x)=x/(1+t)$, the solution also has a closed form: $u(x,t)=\frac{\sum x_{i}^{2}}{2(1+t)}$.. Setting $a=T=2$, we compare Time-DLT with the Deep Galerkin Method (DGM), SirignanoAndSpiliopoulos2018, both implemented using the same ResNet architectures with two residual blocks. DGM minimizes the residual of the PDE on $C$ directly, while our method aims to approximate the analytical solution structure via convex duality. Our experiments, presented in Table 5, show that Time-DLT significantly outperforms DGM in approximating the true solution across various dimensions and times.
Table 5: Comparison of DGM and Time-DLT across dimensions and time horizons.
@@
L2 Error PDE Res. Time (s) L2 Error PDE Res. Time (s) $\rho$ $\sigma$
@@
Near $t=0$, DGM performs better by directly enforcing initial conditions. However, for larger $t$, Time-DLT provides solutions closer to the analytical one [^10] We repeat the experiment 10 times, to get statistically significant values, and report estimated std. deviation $\sigma$.: DGM/DLT L2 error ratio $\rho>1$. This reveals a key insight: while DGM achieves low PDE residuals, minimizing residuals does not necessarily yield the best approximation of the true solution.
@@ 5.5 Symbolic regression with KANs
Kolmogorov–Arnold networks (KANs) have been introduced by KAN. Inspired by the Kolmogorov–Arnold representation theorem, KANs hold great potential in approximation tasks requiring high precision. Moreover, in cases, where the Legendre transform has a closed form expression in terms of special functions, KANs can be employed to obtain the exact solution using symbolic regression. This is a significant advantage compared to numerical approximation since it uncovers the underlying mathematical structure. To illustrate symbolic regression with KANs, we consider the two-dimensional convex functions described in Table 1 above.
In all three cases we performed a symbolic regression with KANs to minimize the empirical squared error (2.2) for a training set $\mathcal{X}_{\rm train}$ sampled uniformly from the square $1,10]^{2}$. We added polynomials up to the third order, $\exp$ and $\log$ as basis functions. The results are given in Table 6. As can be seen, KANs are able to find the correct symbolic expressions for the Legendre transforms of these functions with the right constants (up to rounding errors).
$C$ $D$ residual symbolic expression
quadratic $1,10]^{2} [$$2,20]^{2}$ 1.54e-14 $0.25x_{1}^{2}+0.25x_{2}^{2}$
neg. log $1,10]^{2} [$$1,-0.1]^{2}$ 2.94e-15 $1.0\log(-0.70x_{1})-1.0\log(-6.09x_{2})-0.56$
neg. entropy $1,10]^{2} [$$1,3.3]^{2}$ 4.60e-09 $0.37\exp(1.0x_{1})+0.37\exp(1.0x_{2})$
Table 6: DLT with KANs: two-dimensional training regions, residuals and obtained symb. expressions
@@ 6 Conclusion, limitation and future work
In this paper, we have introduced the Deep Legendre Transform (DLT), a deep learning method for computing convex conjugates of differentiable convex functions $f\colon C\to\mathbb{R}$ defined on an open convex set $C\subseteq\mathbb{R}^{d}$. Our approach provides a posteriori approximation error estimates and scales effectively to high-dimensional settings, outperforming traditional numerical methods.
The limitation of the DLT is that it needs a function $f\colon C\to\mathbb{R}$ to be differentiable and convex. Moreover, it only approximates the convex conjugate $f^{*}$, which can be defined on all of $\mathbb{R}^{d}$, on the image $D=\nabla f(C)$, which is contained in, but in general not equal to, ${\rm dom}\,f^{*}=\left\{y\in\mathbb{R}^{d}:f^{*}(y)<\infty\right\}$.
@@ Appendix A Unbiased a posteriori error estimator
Let $f:C\to\mathbb{R}$ be differentiable and convex on an open convex set $C\subseteq\mathbb{R}^{d}$, and let $D=\nabla f(C)$. Fix a probability measure $\mu$ on $C$ and set $\nu:=\mu\circ(\nabla f)^{-1}$ on $D$. For a measurable $g:D\to\mathbb{R}$, assume
$g-f^{*}\in L^{2}(D,\nu)\quad\text{and}\quad g-f^{*}\in L^{4}(D,\nu)\ \text{(for a finite variance).}$
Let $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu$ and define
$\widehat{\mathcal{E}}_{n}(g)\;:=\;\frac{1}{n}\sum_{i=1}^{n}\Big(g(\nabla f(X_{i}))+f(X_{i})-\langle X_{i},\nabla f(X_{i})\rangle\Big)^{2}.$
Then:
$\text{\emph{(Unbiasedness)}}\qquad\mathbb{E}\big[\widehat{\mathcal{E}}_{n}(g)\big]$ $=\int_{D}\big(g(y)-f^{*}(y)\big)^{2}\,\nu(dy)\;=\;\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}.$
$\text{\emph{(Variance)}}\qquad\mathrm{Var}\big(\widehat{\mathcal{E}}_{n}(g)\big)$ $=\frac{\sigma^{2}}{n},\quad\text{where }\;\sigma^{2}:=\mathrm{Var}\Big(\big(g(\nabla f(X_{1}))-f^{*}(\nabla f(X_{1}))\big)^{2}\Big).$
In particular, by the law of large numbers $\widehat{\mathcal{E}}_{n}(g)\to\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}$ almost surely as $n\to\infty$, and by the central limit theorem $\sqrt{n}\big(\widehat{\mathcal{E}}_{n}(g)-\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}\big)\Rightarrow\mathcal{N}(0,\sigma^{2})$.
@@
By the Fenchel–Young equality $f^{*}(\nabla f(x))=\langle x,\nabla f(x)\rangle-f(x)$, for each $x\in C$ we have
$\Big(g(\nabla f(x))+f(x)-\langle x,\nabla f(x)\rangle\Big)^{2}=\big(g(\nabla f(x))-f^{*}(\nabla f(x))\big)^{2}.$
Taking expectations and using the push-forward measure $\nu=\mu\circ(\nabla f)^{-1}$ yields
$\mathbb{E}\big[\widehat{\mathcal{E}}_{n}(g)\big]=\int_{C}\big(g(\nabla f(x))-f^{*}(\nabla f(x))\big)^{2}\,\mu(dx)=\int_{D}\big(g(y)-f^{*}(y)\big)^{2}\,\nu(dy).$
For the variance, set $Z_{i}:=\big(g(\nabla f(X_{i}))-f^{*}(\nabla f(X_{i}))\big)^{2}$. Then $\widehat{\mathcal{E}}_{n}(g)=\tfrac{1}{n}\sum_{i=1}^{n}Z_{i}$ with i.i.d. $Z_{i}$ and $\mathrm{Var}(\widehat{\mathcal{E}}_{n}(g))=\mathrm{Var}(Z_{1})/n=\sigma^{2}/n$ under the stated moment assumption. The LLN and CLT follow from standard results for i.i.d. averages. ∎

### v2

@@ 5.1 Numerical approximations with MLPs, ResNets and ICNNs
Table 2 shows that DLT matches the performance of direct learning both in terms of approximation accuracy and training time. This empirical finding is maintained across different network architectures and dimensions, confirming that the performance of DLT is only constrained by the neural networks’ capacity to approximate multidimensional functions and not by its implicit formulation. In our experiments, ResNet provided the best approximations despite not guaranteeing convexity, while MLP-ICNN showed limitations due to the lack of skip connections compared to ICNN [^9] Additional multidimensional experiments confirming that DLT has the same performance as direct learning of known conjugates in terms of approximation accuracy and training time are given in the Appendix..
Table 2: Comparison of DLT with direct learning (Dir) for $d=10$ and $50$. Errors were computed with Monte Carlo on test set of $4096$ i.i.d. random points sampled from $\text{Unif}\circ(\nabla f)^{-1}$.
@@
$_{\text{DLT}} / RMSE$$_{\text{Dir}}$ $t_{\rm train}$ (s) $_{\text{DLT}} / RMSE$$_{\text{Dir}}$ $t_{\rm train}$ (s)
@@
Functions without closed form conjugate
Most functions do not admit a closed form expression for the convex conjugate. However, DLT gives the same a posteriori error estimates as if $f^{*}$ were known explicitly and learned directly on $D$. For illustration, we consider the quadratic over linear function
$f(x)=\frac{x^{2}_{1}+\dots+x^{2}_{d}+1}{x_{1}+\dots+x_{d}+1}.$
It is not defined everywhere on $\mathbb{R}^{d}$. But it can be shown that it is convex on the convex set $C=\{x\in\mathbb{R}^{d}:x_{1}+\dots+x_{d}>0\}$; see e.g. Boyd and Vandenberghe (2004). The $i$-th component of the gradient $\nabla f(x)$ is
@@
But the convex conjugate $f^{*}$ is not known in closed form, and the image $D=\nabla f(C)$ is a non-trivial subset of $\mathbb{R}^{d}$. Even though we do not know $f^{*}$ explicitly, DLT yields accurate ICNN approximations of $f^{*}$ on $D$. Here, we used an ICNN with 3 hidden layers and ELU activations, we sampled random points from a standard half-normal distribution on $C$. Figure 1 depicts these random points and their images under the gradient map $\nabla f$ given in (5.1) for $d=2$. The table on the right shows estimates of the $L^{2}$-approximation error $\left\|g_{\hat{\theta}}-f^{*}\right\|_{L^{2}(\mathbb{R}^{d},\nu)}$ together with the times in minutes and seconds it took to train the approximations $g_{\hat{\theta}}$ for different $d$. Note that the true $f^{*}$ is unknown here, but due to the results of Section 3, it is still possible to provide $L^{2}$ approximation errors.
In Appendix B, we study the following additional non-separable examples: (i) quadratic $f(x)=\frac{1}{2}x^{\top}Qx$; (ii) exponential-minus-linear $f(x)=e^{\langle a,x\rangle}-\langle b,x\rangle$; and (iii) ICNN-represented functions.
$d$ $\|g_{\theta}-f^{*}\|_{L^{2}}$ Time ($m$:$s$)
@@ 5.2 Comparison with classical grid methods
Next, we compare DLT against the LLT algorithm of Lucet (1997), a state-of-the-art grid-based method for computing convex conjugates. Rather than directly evaluating $\max_{x\in\mathcal{X}}\{\langle x,y\rangle-f(x)\}$ over a full $d$-dimensional grid $\mathcal{X}$ of size $N^{d}$ (which requires $O(N^{2d})$ operations), Lucet reformulates the $d$-dimensional maximization as a sequence of nested 1D problems:
$f^{*}(y_{1},\ldots,y_{d})=\max_{x_{1}}\left\{y_{1}x_{1}+\max_{x_{2}}\left\{y_{2}x_{2}+\cdots+\max_{x_{d}}\left\{y_{d}x_{d}-f(x)\right\}\right\}\right\}.$
This decomposition reduces the computational complexity to $O(dN^{d+1})$ while still requiring $O(N^{d})$ memory. Table 3 compares the methods for a fixed discretization of $N=10$ points per dimension for two of the functions from Table 1. All errors were evaluated on the same random sample $\mathcal{Y}_{\rm test}\subseteq D=\nabla f(C)$ against the true $f^{*}$. For low dimensions ($d\in\{2,6\}$), Lucet runs quickly with modest memory, making it possible to use much finer grids within a reasonable time budget. However, as $d$ grows beyond 88, even modest grids become memory or time infeasible due to $O(N^{d})$ storage and $O(d\cdot N^{d+1})$ computational complexity. In contrast, DLT maintains a constant memory footprint and scales smoothly to dimensions $d$ beyond 9.
Table 3: Comparison of Lucet with DLT for dimensions $d\in\{2,6,8,10\}$. Lucet uses $N=10$ points per dim; DLT uses a ResNet(128,128)(128,128). Reported: solution time $t_{\mathrm{solve}}$, evaluation time $t_{\mathrm{eval}}$, active memory (MB) and RMSE.
@@
When the gradient mapping $\nabla f$ is significantly distorting, the image $\nabla f(\mathcal{X})$ of a uniformly sampled training set $\mathcal{X}\subseteq C$ will be highly non-uniformly distributed on $D$. This skewed distribution can adversely affect learning dynamics, concentrating approximation accuracy in certain regions while neglecting others. The approximate inverse sampling method introduced in Section 4 addresses this issue by allowing us to sample from any desired distribution (e.g., uniform) directly on $D$.
@@
Table 4: Comparison of direct and inverse sampling for the negative logarithm. Inverse sampling uses an approximate inverse map $h_{\hat{v}}$ whose quality is measured by $2/(s\sqrt{d})\|\nabla f\circ h_{\hat{v}}-\text{id}\|_{L^{1}}$, $s=990$ is the side-length of $D$ and total training time split between $h_{\hat{v}}$ and $g_{\hat{\theta}}$.
@@ 5.4 Applications to Hamilton–Jacobi equations
The Hamilton–Jacobi (HJ) equation is a cornerstone of classical mechanics, optimal control theory, and mathematical physics, which provides a natural application domain for DLT.
@@
for two convex functions $H,g:\mathbb{R}^{d}\to\mathbb{R}$.
The Hopf formula (Hopf, 1965) provides the following expression of $u$ as a double convex conjugate:
@@
While this formula is explicit, evaluating it in high dimensions remains computationally challenging. To address this, we propose Time-DLT (Time-Parameterized Deep Legendre Transform) which approximates $(g^{*}+tH)^{*}$ using DLT with a neural network $u_{\theta}(x,t)$ with input variables $x$ and $t$.
@@
We study an HJ equation (5.1) with $H(x)=g(x)=\frac{1}{2}\sum_{i=1}^{d}x^{2}_{i}$. It has the explicit solution $u(x,t)=\frac{\sum_{i=1}^{d}x_{i}^{2}}{2(1+t)}$, which we can use a reference. We use Time-DLT to compute a solution on $C=(-a,a)^{d}\times[0,T]$ for $a=T=2$ and compare against the DGM (Deep Galerkin Method) of Sirignano and Spiliopoulos (2018), both implemented using the same ResNet architectures with two residual blocks. DGM minimizes the residual of the PDE on $C$ directly, while Time-DLT approximates the Hopf formula (5.2). Table 5 shows that Time-DLT outperforms DGM in approximating the true solution across different dimensions and times (however, we used the explicit expression for $g^{*}$ and just computed the outer Legendre transform).
Table 5: Comparison of DGM and Time-DLT for different dimensions and times.
@@
$L^{2}$ Error PDE Res. time (s) $L^{2}$ Error PDE Res. time (s) $\rho$ $\sigma$
@@
Near $t=0$, DGM performs better by directly enforcing the initial condition. But for larger $t$, the Time-DLT solutions are closer to the correct ones [^10] We repeat the approximations 10 times to get statistically significant results and report estimated std. deviations $\sigma$. indicated by a DGM/DLT $L^{2}$ error ratio $\rho>1$. Note that DGM aims to minimize PDE residuals, but small PDE residuals do not always correspond to small approximation errors.
@@ 5.5 Symbolic regression with KANs
Kolmogorov–Arnold networks (KANs) have recently been introduced by Liu et al. (2025). Inspired by the Kolmogorov–Arnold representation theorem, KANs hold great potential in approximation tasks requiring high precision. Moreover, in cases, where the Legendre transform has a closed form expression in terms of special functions, KANs can be employed to obtain the exact solution using symbolic regression. This is a significant advantage compared to numerical approximation since it uncovers the underlying mathematical structure. To illustrate symbolic regression with KANs, we consider the three functions of Table 1 in two dimensions.
In all three cases we performed a symbolic regression with KANs to minimize the empirical squared error (2.2) for a training set $\mathcal{X}_{\rm train}$ sampled uniformly from the square $1,10]^{2}$. We used polynomials up to the third order, $\exp$ and $\log$ as basis functions. As can be seen from the results in Table 6, KANs are able to find the correct symbolic expressions for the Legendre transforms of these functions with the right constants (up to rounding errors).
Table 6: DLT with KANs: two-dimensional training regions, $L^{2}$-errors and obtained symbolic expressions.
Function $C$ $D$ $L^{2}$ Error Symbolic Expression
$\sum_{i=1}^{2}x^{2}_{i}$ $1,10]^{2} [$$2,20]^{2}$ 0 $0.25y_{1}^{2}+0.25y_{2}^{2}$
$\sum_{i=1}^{2}\log(x_{i})$ $1,10]^{2} [$$1,-0.1]^{2}$ 5.42e-8 $1.0\log(-0.70y_{1})-1.0\log(-6.09y_{2})-0.56$
$\sum_{i=1}^{2}x_{i}\log(x_{i})$ $1,10]^{2} [$$1,3.3]^{2}$ 6.78e-05 $0.37\exp(1.0y_{1})+0.37\exp(1.0y_{2})$
@@ 6 Conclusion, limitation and future work
In this paper, we have introduced DLT (Deep Legendre Transform), a deep learning method for computing convex conjugates of differentiable convex functions $f\colon C\to\mathbb{R}$ defined on an open convex set $C\subseteq\mathbb{R}^{d}$. Our approach provides a posteriori approximation error estimates and scales effectively to high-dimensional settings, outperforming traditional numerical methods.
The limitation of DLT is that it needs a function $f\colon C\to\mathbb{R}$ to be differentiable and convex. Moreover, it only approximates the convex conjugate $f^{*}$, which can be defined on all of $\mathbb{R}^{d}$, on the image $D=\nabla f(C)$, which is contained in, but in general not equal to, ${\rm dom}\,f^{*}=\left\{y\in\mathbb{R}^{d}:f^{*}(y)<\infty\right\}$.
@@ Appendix A Unbiased a posteriori error estimator
Let $f\colon C\to\mathbb{R}$ be a differentiable convex function on an open convex set $C\subseteq\mathbb{R}^{d}$ and denote $D=\nabla f(C)$. Fix a probability measure $\mu$ on $C$ and consider the push-forward $\nu=\mu\circ(\nabla f)^{-1}$ on $D$. For a measurable function $g\colon D\to\mathbb{R}$, denote
$\widehat{\mathcal{E}}_{n}(g)=\frac{1}{n}\sum_{i=1}^{n}\Big(g(\nabla f(X_{i}))+f(X_{i})-\langle X_{i},\nabla f(X_{i})\rangle\Big)^{2},$
where $X_{1},\dots,X_{n}$ are i.i.d. random vectors with distribution $\mu$. Then the following hold:
  • (i)
    If $g-f^{*}\in L^{2}(D,\nu)$, one has
    $\mathbb{E}\!\left[\widehat{\mathcal{E}}_{n}(g)\right]=\int_{D}\Big(g(y)-f^{*}(y)\Big)^{2}\,\nu(dy)=\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}\quad\mbox{(Unbiasedness)}$
    and $\widehat{\mathcal{E}}_{n}(g)\to\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}$ almost surely for $n\to\infty$.
  • (ii)
    If $g-f^{*}\in L^{4}(D,\nu)$, then
    $\mathrm{Var}\left(\widehat{\mathcal{E}}_{n}(g)\right)=\frac{\sigma^{2}}{n}\quad\text{for}\quad\sigma^{2}=\mathrm{Var}\left(\left(g(\nabla f(X_{1}))+f(X_{1})-\left<X_{1},\nabla f(X_{1})\right>\right)^{2}\right)$
    and $\sqrt{n}\left(\widehat{\mathcal{E}}_{n}(g)-\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}\right)\to\mathcal{N}(0,\sigma^{2})$ in distribution as $n\to\infty$. In particular, if $\widehat{\sigma}$ is an estimate of $\sigma$ and $z_{\alpha}$ denotes the $\alpha$-quantile of $\mathcal{N}(0,1)$, then
    $\widehat{\mathcal{E}}_{n}(g)\pm\frac{\widehat{\sigma}}{\sqrt{n}}z_{1-\alpha/2}$
    is an approximate $(1-\alpha)$-confidence interval for $\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}$.
@@
(i) By the Fenchel–Young equality $f^{*}(\nabla f(x))=\langle x,\nabla f(x)\rangle-f(x)$, one has
$\Big(g(\nabla f(X_{i}))+f(X_{i})-\langle X_{i},\nabla f(X_{i})\rangle\Big)^{2}=\big(g(\nabla f(X_{i}))-f^{*}(\nabla f(X_{i}))\big)^{2}.$
So, taking expectation and using the push-forward measure $\nu=\mu\circ(\nabla f)^{-1}$ yields
$\mathbb{E}\!\left[\widehat{\mathcal{E}}_{n}(g)\right]\hskip-2.84526pt=\hskip-2.84526pt\int_{C}\big(g(\nabla f(x))-f^{*}(\nabla f(x))\big)^{2}\,\mu(dx)\hskip-0.56905pt=\hskip-2.84526pt\int_{D}\big(g(y)-f^{*}(y)\big)^{2}\,\nu(dy)\hskip-0.85358pt=\hskip-0.85358pt\|g-f^{*}\|^{2}_{L^{2}(D,\nu)},$
and it follows from the law of large numbers that $\widehat{\mathcal{E}}_{n}(g)\to\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}$ almost surely. This shows (i).
(ii) Denote
$Z_{i}=\Big(g(\nabla f(X_{i}))+f(X_{i})-\left<X_{i},\nabla f(X_{i})\right>\Big)^{2}.$
Then $\widehat{\mathcal{E}}_{n}(g)=\tfrac{1}{n}\sum_{i=1}^{n}Z_{i}$, and therefore,
$\mathrm{Var}\left(\widehat{\mathcal{E}}_{n}(g)\right)=\frac{\mathrm{Var}(Z_{1})}{n}=\frac{\sigma^{2}}{n}.$
So, it follows from the central limit theorem that $\sqrt{n}\left(\widehat{\mathcal{E}}_{n}(g)-\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}\right)\to\mathcal{N}(0,\sigma^{2})$ in distribution. In particular, $\widehat{\mathcal{E}}_{n}(g)\pm\frac{\widehat{\sigma}}{\sqrt{n}}z_{1-\alpha/2}$ is an approximate $(1-\alpha)$-confidence interval for $\|g-f^{*}\|^{2}_{L^{2}(D,\nu)}$, which concludes the proof. ∎